OpenStudy (anonymous):

one order of 3 burritos and 4 tacos which came up to 11.33 and the other person ordered 9 burritos and 5 tacos cost 23.56.. what is the cost of the taco

OpenStudy (anonymous):

burriro =x taco = y 3x + 4y = 11.33 9x+5y = 23.56

OpenStudy (anonymous):

it said the cost of 1 taco

OpenStudy (anonymous):

@nphuongsun93

OpenStudy (anonymous):

Typo, Let the Mexican Quest begin. (Note: I'm Hispanic, the joke stands). So let the price of your tacos a variable, say \(x\), and the price of burritos be the other variable, say \(y\). So, we need two equations to solve for two variables. One of them will be the cost of the first person and the other one will be the cost of the second person, so: \(3Burritos+4Tacos=$11.33\) or \(3y+4x=11.33\) \(9Burritos+5Tacos=$23.56\) or \(9y+5x=23.56\) Solve the system of for x and that will be the price of one taco.

OpenStudy (anonymous):

@jlvm you messed up the variables D: ..

OpenStudy (anonymous):

I noticed it before you even wrote it, but thanks.

OpenStudy (anonymous):

3x + 4y = 11.33 9x + 5y = 23.56 3x + 4y = 11.33 times 3 = 9x + 12y = 33.99 9x + 12y = 33.99 - 9x + 5y = 23.56 ----------------------- 7y = 10.43 y = 1.49 now sub y into the one of the equations to find x... actually you don't even need to, you only need to find y, the cost of taco

OpenStudy (anonymous):

Now I want a taco so bad...@davonnn you owe me a taco.

OpenStudy (alexwee123):

tacos!! :D